A box with a volume V = 0.0500 cubic meter lies at the bottom of a lake whose water has a density of 1.0 X 103kg/cubic meter. How much force is required to lift the box, if the box has a mass of a. 1,000 kg, b. 100 kg and c. 550 kg?

1
Expert's answer
2021-02-19T18:58:19-0500

Given,

Volume of water, "V=0.05m^3"

density, "\rho=1000kg/m^3"

Accelaration due to gravity, "g=9.8m\/s^2"


The buoyant force on the box is

          "=V\times \rho\times g 

 \\=0.05\times 1000\times 9.8

 \\ = 490 \text{ N}"


Case-1: When box mass is 1000kg

Then force require to lift

"F_{\text{ lift}}=mg-\text{Buoyant force}"

"=1000(9.8)-490\\=9800-490\\=9310N"



Case-2: When box mass is 100kg

Then force require to lift

"F_{\text{ lift}}=mg-\text{Buoyant force}"

"=100(9.8)-490\\=980-490\\=490N"



Case-3: When box mass is 550kg

Then force require to lift

"F_{\text{ lift}}=mg-\text{Buoyant force}"

"=550(9.8)-490\\=5390-490\\=4900N"